OpenStudy (anonymous):

Find the area of a square chessboard that measures 15 inches on each side A. 225 in² B. 60 in² C. 30 in² D. 15 in²

OpenStudy (gorv):

area= side*side

OpenStudy (gorv):

side=15

OpenStudy (anonymous):

ok

OpenStudy (gorv):

plug in 15 in formula

OpenStudy (anonymous):

it's 15in2d

OpenStudy (anonymous):

is it d

OpenStudy (anonymous):

is it d

OpenStudy (anonymous):

What's 15 x 15?

OpenStudy (idealist10):

The answer is A. Because A=s^2 where side=15 inches. So 15*15=225.

OpenStudy (anonymous):

225

OpenStudy (anonymous):

ok it's a

OpenStudy (anonymous):

thank you

OpenStudy (idealist10):

You're welcome.

OpenStudy (anonymous):